# Information Theory for Complex Systems Scientists: What, Why, & How?

### v3

@@ VI.4 Do We Even Need Complexity Measures?
The question of a scalar measure of complexity is a naturally interesting one, and, as mentioned, considerable work has gone into attempting to derive the most “natural” one. It is worth taking a step back and, borrowing from the title of [226] ask: ”measures of statistical complexity: why?” As mathematical question, the project of finding a scalar measure that captures the nuances of what it means to be complex is an interesting one, but ultimately it is often unclear what has been learned by announcing that “this system has x units of complexity”. While there are some cases where relative changes in complexity can be informative (for example, see [227, 228, 229, 230, 142, 143] for examples), in general it is more useful to begin by defining what particular process or dynamic one is trying to quantify and going from there. Is it synergy/redundancy dominance? Memory? Integration? Predictability? As complex systems science continues to mature, the ability to answer specific questions about specific aspects of system dynamics will become increasingly important, while vaguer, more abstract notions about “complexity” may have to fall by the wayside.
@@ VII Information Estimators
In this section, we will discuss how to estimate entropies and mutual informations for discrete and continuous data. The discrete case is reasonably straightforward (although, as discussed in Sec. V.1, bias associated with finite size effects is a concern). The situation becomes more complex for continuous data. We will discuss parametric (Gaussian) estimators for continuous data, as well as briefly touching on non-parametric estimators using K-nearest neighbor graphs. Each one of these has its own strengths and weaknesses, and critical considerations should be made at the outset of an analysis about what approach is best. This section will not cover all possible families of information estimators: for example, there is a small, but growing field concerned with the problem of estimating information-theoretic dependencies between point processes in continuous time [231, 232, 233, 234], with application to processes such as spiking neurons and heart-rate. Similarly, the problem of estimating mutual information between mixtures of discrete and continuous variables is an area of research under active development [235, 236].To those interested in these more niche applications, we refer to the aforementioned literature.
@@ VII.1 Discrete Signals
This is variously known as the naive entropy estimator, the plugin estimator, as well as the maximum likelihood estimator, and is probably the most commonly used method for estimating entropies from data. Despite its frequency of use, however, the plugin estimator has fundamental limitations: the most significant of which is that it systematically under-estimates the true entropy of the process. This is known as bias ($B$) and is given by:
$\text{E}[\hat{H}(X)]\leq H(X)$ (104)
The severity of the bias is a function of both the number of states of the system and the number of samples used to construct the probability distributions [172], and a bias-correction term can be introduced for a bias-corrected estimator:
@@
The mutual information suffers from a similar bias, although unlike the entropy, the plugin estimator of mutual information will generally over-estimate the true information quantity. Since the mutual information is often calculated as a difference in multiple entropy measures (eg. Eqs. 26), if the bias in MLE estimates of the entropy is greater than the difference, significant errors can be introduced. A number of bias-corrected estimators of mutual information exist (see: [237]), although the majority of work has been done on estimating the mutual information from continuous signals, although the risk of bias is also of concern even when looking at naturally discrete data.
@@ VII.2.1 Coarse Graining
An alternative to binning the data is creating a “point-process”, which maps the continuous signal to a binary time-series by marking “events” at the most extreme points of the signal, usually above some threshold. Point processing assumes that only a subset of moments (typically unusually high-amplitude samples) are significant or informative about the underlying dynamics, and that the majority of the signal is essentially noise. Assuming the assumptions are valid, point-processing is a much more natural method of compressing a complex, continuous signal, as it only records significant events and does not conflate signal and noise in the same way that histogram binning does.
@@
Both the histogram-based binning and the point-processing transformation assume that every observation of a time-series is a random draw from some underlying distribution, ignoring the possibility that temporally-extended patterns are meaningful as well. An alternative coarse-graining approach is the permutation embedding or ordinal partition embedding, which transforms continuous data into a sequence based on temporally-extended amplitude patterns [241].
@@ VII.2.2 Differential Entropy
This satisfies the definition that $H^{dx}(X)=\mathbb{E}[-\log P(x)]$, however, it fails to preserve several other properties that make the discrete entropy a useful measure [44]. For instance, it can be negative (as in the case of very tight Gaussian distributions), and it is not invariant under under rescaling due to the additive nature of the logarithm:
@@
While the differential entropy is not entirely well-behaved, the differential Kullback-Leibler divergence maintains the desire properties of the discrete divergence:
@@ VII.2.3 Gaussian Estimators
Where $\sigma^{2}$ is the standard deviation and $|\Sigma|$ is the determinate of the $N$-dimensional covariance matrix. For a continuous, univariate distribution on the interval [,][$\infty$,$\infty$][ - ∞ , ∞ ] and with fixed mean $\mu$ and variance $\sigma$, the Gaussian distribution is the maximum entropy distribution. For a multivariate distribution with infinite support, fixed means, and fixed, pairwise covariances, the multivariate Gaussian is the maximum entropy distribution possible [44].
@@
$I^{dx}(X;Y)$ $=\frac{-\ln(1-\rho^{2})}{2}$ (111)
Where $\rho$ is the Pearson correlation coefficient between $X$ and $Y$. If X and Y are multivariate, the generalized Gaussian mutual information is of the form:
@@
$P(x)=\frac{1}{\sigma\sqrt{2\partial}}e^{-\frac{1}{2}\left(\frac{x-\mu}{\sigma}\right)^{2}}$ (114)
@@
$P(x_{1},x_{2}...x_{n})=\frac{1}{(2\partial)^{\frac{n}{2}}|\Sigma|^{\frac{1}{2}}}e^{\frac{-(\boldsymbol{x-\mu})^{T}\Sigma^{-1}(\boldsymbol{x-\mu})}{2}}$ (115)
@@ VII.2.4 Density-Based Estimators Estimators
With any KNN-based analysis, it is crucial to pre-specify the particular distance function to be used and assess its appropriateness. While in principle any valid metric could be used for the following estimators, in practice, concerns such as the curse of dimensionality often limit what measures will be useful. Here, we are assuming that the reader is planning to use the $L_{i}nfty$ (Chebyshev) distance, which is robust to high dimensional data and simplifies some of the equations. Readers interested in alternatives such as the $L_{1}$ (Manhattan distance) or the $L_{2}$ (Euclidean distance) should review the cited literature.
@@ X Conclusions
In this piece, we have attempted to provide a robust (although far from comprehensive) introduction to the theory and applications of multivariate information theory to the study of complex systems. In doing so, we hope to make information theory a more widely known and commonly used tool in the toolkit of modern science. We feel that this document should not be the only exposure to information theory that an interested reader explores: different authors present the same material in different ways, and there is value in viewing things from different perspectives. Textbooks such as Bossomaier et al., [84], Cover and Thomas [44], and McKay [26] were all consulted during the writing of this review and are themselves invaluable resources. There are also topics that this review does not even touch on, including maximum entropy models, source coding, information bottlenecks, the links between information entropy and physical entropy, and the role of information theory in machine learning. While these are all deep topics worth exploration, they are either beyond the scope of this review (such as information theory in machine learning) or require concepts from other domains that may not be relevant to complex systems scientists as a whole (such as links to thermodynamics or maximum-entropy models).

### v4

@@ VI.4 Do We Even Need Complexity Measures?
The question of a scalar measure of complexity is a naturally interesting one, and, as mentioned, considerable work has gone into attempting to derive the most “natural” one. It is worth taking a step back and, borrowing from the title of [230] ask: ”measures of statistical complexity: why?” As a mathematical question, the project of finding a scalar measure that captures the nuances of what it means to be complex is an interesting one, but ultimately it is often unclear what has been learned by announcing that “this system has x units of complexity”. While there are some cases where relative changes in complexity can be informative (for example, see [231, 232, 233, 234, 144, 145] for examples), in general it is more useful to begin by defining what particular process or dynamic one is trying to quantify and going from there. Is it synergy/redundancy dominance? Memory? Integration? Predictability? As complex systems science continues to mature, the ability to answer specific questions about specific aspects of system dynamics will become increasingly important, while vaguer, more abstract notions about “complexity” may have to fall by the wayside.
@@ VII Information Estimators
In this section, we will discuss how to estimate entropies and mutual informations for discrete and continuous data. The discrete case is reasonably straightforward (although, as discussed in Sec. V.1, bias associated with finite size effects is a concern). The situation becomes more complex for continuous data. We will discuss parametric (Gaussian) estimators for continuous data, as well as briefly touching on non-parametric estimators using K-nearest neighbor graphs. Each one of these has its own strengths and weaknesses, and critical considerations should be made at the outset of an analysis about what approach is best. This section will not cover all possible families of information estimators: for example, there is a small, but growing field concerned with the problem of estimating information-theoretic dependencies between point processes in continuous time [235, 236, 237, 238], with application to processes such as spiking neurons and heart-rate. Similarly, the problem of estimating mutual information between mixtures of discrete and continuous variables is an area of research under active development [239, 240]. Finally, Bayesian estimators have been developed with many appealing properties [241]. To those interested in these more niche applications, we refer to the aforementioned literature.
@@ VII.1 Discrete Signals
This is variously known as the naive entropy estimator, the plugin estimator, as well as the maximum likelihood estimator, and is probably the most commonly used method for estimating entropies from data. Despite its frequency of use, however, the plugin estimator has fundamental limitations: the most significant of which is that it systematically under-estimates the true entropy of the process. This is known as bias and is given by:
$\mathbb{E}[\hat{H}(X)]\leq H(X)$ (104)
The severity of the bias is a function of both the number of states of the system and the number of samples used to construct the probability distributions [175], and a correction term can be introduced for a bias-corrected estimator:
@@
The mutual information suffers from a similar bias, although unlike the entropy, the plugin estimator of mutual information will generally over-estimate the true information quantity. Since the mutual information is often calculated as a difference in multiple entropy measures (eg. Eqs. 26), if the bias in MLE estimates of the entropy is greater than the difference, significant errors can be introduced. A number of bias-corrected estimators of mutual information exist (see: [242]), although the majority of work has been done on estimating the mutual information from continuous signals, although the risk of bias is also of concern even when looking at naturally discrete data. Finally, the bias in entropy can also impact discrete estimators of higher-order measures such as the O-information, biasing it towards synergy and these measures can require correction when the number of states is much larger than the number of samples [243].
@@ VII.2.1 Coarse Graining
An alternative to binning the data that applies naturally to time series data is creating a “point-process”, which maps the continuous signal to a binary time-series by marking “events” at the most extreme points of the signal, usually above some threshold. Point processing assumes that only a subset of moments (typically unusually high-amplitude samples) are significant or informative about the underlying dynamics, and that the majority of the signal is essentially noise. Assuming the assumptions are valid, point-processing is a much more natural method of compressing a complex, continuous signal, as it only records significant events and does not conflate signal and noise in the same way that histogram binning does. When analyzing point-processed data, one can use a standard discrete entropy estimator, or more recent and sophisticated measures of continuous-time entropy that operate on inter-even intervals rather than treating each frame as a draw from a fixed distribution [236, 237, 238].
@@
for time series data, the histogram-based binning transformation assumes that every observation of a time-series is a random draw from some underlying distribution, ignoring the possibility that temporally-extended patterns are meaningful as well. An alternative coarse-graining approach is the permutation embedding or ordinal partition embedding, which transforms continuous data into a sequence based on temporally-extended amplitude patterns [247].
@@ VII.2.2 Differential Entropy
This satisfies the definition that $H^{dx}(X)=\mathbb{E}[-\log(P(x))]$, however, it fails to preserve several other properties that make the discrete entropy a useful measure [43]. For instance, it can be negative (as in the case of very tight Gaussian distributions), and it is not invariant under under rescaling due to the additive nature of the logarithm:
@@
While the differential entropy is not entirely well-behaved, the differential Kullback-Leibler divergence maintains the desired properties of the discrete divergence:
@@ VII.2.3 Gaussian Estimators
Where $\sigma$ is the variance and $|\Sigma|$ is the determinate of the $N$-dimensional covariance matrix. For a continuous, univariate distribution on the interval [,][$\infty$,$\infty$][ - ∞ , ∞ ] and with fixed mean $\mu$ and variance $\sigma$, the Gaussian distribution is the maximum entropy distribution. For a multivariate distribution with infinite support, fixed means, and fixed, pairwise covariances, the multivariate Gaussian is the maximum entropy distribution possible [43].
@@
$I^{dx}(X;Y)$ $=\frac{-\ln(1-r^{2})}{2}$ (111)
Where $r$ is the Pearson correlation coefficient between $X$ and $Y$. If X and Y are multivariate, the generalized Gaussian mutual information is of the form:
@@
$P(x)=\frac{1}{\sigma\sqrt{2\sigma^{2}}}e^{-\frac{1}{2}\left(\frac{x-\mu}{\sigma^{2}}\right)^{2}}$ (114)
@@
$P(x_{1},x_{2}...x_{n})=\frac{1}{(2\pi)^{\frac{n}{2}}|\Sigma|^{\frac{1}{2}}}e^{\frac{-(\boldsymbol{x-\mu})^{T}\Sigma^{-1}(\boldsymbol{x-\mu})}{2}}$ (115)
@@ VII.2.4 Density-Based Estimators Estimators
With any KNN-based analysis, it is crucial to pre-specify the particular distance function to be used and assess its appropriateness. While in principle any valid metric could be used for the following estimators, in practice, concerns such as the curse of dimensionality often limit what measures will be useful. Here, we are assuming that the reader is planning to use the $L_{\infty}$ (Chebyshev) distance, which is robust to high dimensional data and simplifies some of the equations. Readers interested in alternatives such as the $L_{1}$ (Manhattan distance) or the $L_{2}$ (Euclidean distance) should review the cited literature.
@@ X Conclusions
In this piece, I have attempted to provide a robust (although far from comprehensive) introduction to the theory and applications of multivariate information theory to the study of complex systems. In doing so, I hope to make information theory a more widely known and commonly used tool in the toolkit of modern science. I feel that this document should not be the only exposure to information theory that an interested reader explores: different authors present the same material in different ways, and there is value in viewing things from different perspectives. Textbooks such as Bossomaier et al., [83], Cover and Thomas [43], and McKay [26] were all consulted during the writing of this review and are themselves invaluable resources. There are also topics that this review does not even touch on, including maximum entropy models, source coding, information bottlenecks, the links between information entropy and physical entropy, and the role of information theory in machine learning. While these are all deep topics worth exploration, they are either beyond the scope of this review (such as information theory in machine learning) or require concepts from other domains that may not be relevant to complex systems scientists as a whole (such as links to thermodynamics or maximum-entropy models).